(b) (i)Design a logic circuit that will allow a signal to pass to the output only when control inputs B and C are both HIGH; otherwise, the output will stay LOW. (4marks)CR
(ii) Design a logic circuit that allows a signal to pass to the output only when one, but not both, of the control inputs are HIGH; otherwise, the output will stay HIGH. (4marks)CR

(c) What is a universal gate? Give examples. Realize the basic gates with any one universal gate. (5marks) AN
        (Total 20marks)
Answer #1

and a are Gullen (i) Design logic circuit only when control inputs B both HIGH → allow a signal to pass the output will stayoutpub (%) - ABC ABC =(A+A BC (-:- Ata =D - Вс .: X-BC logic digital circuit:- B х с (ii) Decign alogie Girait conted inpectsJulpub (x) = ĀBCT ĀBit ABCT ABT (B) +A (BOC) -CA79) (BOC) (.. Ata =1) =BctB .:XBC+B7 logic Ertuiti B X (outpull) (c) universa